OpenStudy (colcaps):

What are the degree measures of A and B if C = 54, A = x and B = 2x for triangle ABC?

OpenStudy (ajprincess):

54+x+2x=180 Can u find nw?

OpenStudy (colcaps):

a.15 and 30 b. 18 and 36 c. 21 and 42 d. 25 and 50

hartnn (hartnn):

angles in a triangle sum up to 180 so u get x+2x+54=180 can u solve this to get x ??

OpenStudy (theviper):

x+2x+54=180 3x+54=180 3x=180-54=126 x=126/3=42

OpenStudy (theviper):

A=x=42 B=2x=2*42=84

OpenStudy (anonymous):

<A + <B + <C = 180

OpenStudy (theviper):

Yes by angle sum property of a triangle ;)

OpenStudy (colcaps):

A=X B=2x

OpenStudy (colcaps):

so i have to find that?

OpenStudy (theviper):

Put the value of x that is 42 A=x=42 B=2x=2*42=84

OpenStudy (colcaps):

so A=42 B=21??

OpenStudy (colcaps):

oh

OpenStudy (anonymous):

so 54 + x + 2x = 180 54 + 3x = 180 3x = 180 - 54 = 126 Hence x = 42, 2x = 2(42) = 84 so A =42, B = 84

OpenStudy (theviper):

B=2x= 2*42=84

OpenStudy (theviper):

@Colcaps

OpenStudy (colcaps):

but the answer doesent answer my problem the coices are here 15 and 30 18 and 36 21 and 42 25 and 50

OpenStudy (colcaps):

wait

OpenStudy (colcaps):

21 and 42 is simplified thank you guys

OpenStudy (anonymous):

Then its possible that some of the angles are exterior angles

OpenStudy (anonymous):

I suppose x and 2x to be exterior

OpenStudy (colcaps):

i found the answer its 21 and 42 im not looking for the angle but thanks anyway

OpenStudy (anonymous):

Then x + 2x = 54 3x = 54 x = 18, 2x = 2(18) = 36

OpenStudy (mayankdevnani):

 so